# A Temporal-Spectral Fusion Transformer with Subject-specific Adapter for Enhancing RSVP-BCI Decoding

### v1

@@ 4.3 Fusion Module
The fusion module is proposed after the cross-view interaction module to fuse temporal and spectral tokens, thereby generating comprehensive fusion features for classification (see Fig. 4). Since temporal tokens and spectral tokens are extracted from EEG temporal signals and spectrogram images during the same time period, they naturally align in the time dimension. Initially, we token-wise concatenate the two-view tokens and linearly project them into a $d$-dimensional feature space, each of them containing temporal-spectral information for its corresponding time period. Then we designate these tokens as queries and employ both temporal and spectral tokens as keys and values in a cross-attention layer to obtain fusion tokens ($\boldsymbol{X}_{fus}\in\mathbb{R}^{n\times d}$). At the end of the fusion module, we design a convolution layer with 16 convolution kernels to aggregate all fusion tokens and flatten the output feature into a feature vector as the fusion feature ($\boldsymbol{z}_{fus}$) for final prediction. The convolution kernel size is $(16,d/8)$ with a stride of $(16,d/8)$.
@@ 4.4 Subject-specific Adapter
Owing to variations in data distribution among different subjects, the pre-trained model needs to transfer the learned general patterns to the data distribution of new subjects for efficient decoding. The adapter-based fine-tuning can realize the transfer of pre-trained models to downstream tasks through updating a few parameters, which can avoid overfitting caused by limited data houlsby2019parameter . Thus, we introduce a Subject-specific Adapter (SA) to rapidly transfer the knowledge of pre-trained model learned from existing subjects’ data to data characteristics of new subject using a small amount of data from the new subject.
@@ 5 Experiments
5.1 Implementation Details
We employ the Mexican hat wavelet basis function for the CWT with a scale set to 20, as P300 detection emphasizes low-frequency EEG information. There are four hyperparameters in our model, including the slice length ($t$), the embedding dimension ($d$), the number of heads ($h$) in the multi-head attention mechanism, and the temperature ($\tau$) in multi-view consistency learning. The slice length ($t$) is set to 5 in the slice embedding layer. To prevent overfitting, we appropriately reduce the size of the model by setting $d$ as 128 and $h$ as 4. For multi-view consistency learning, we set the temperature parameter ($\tau$) at 0.2. Furthermore, for simplicity, we equate $d_{v}$ to $d_{k}$ within the attention mechanism.
We implement our model within the PyTorch framework. In both training stages, we utilize the Adam optimizer kingma2014adam for model optimization with a learning rate set to 0.0005, which is reduced by 20% every 40 epochs. Additionally, the L2 regularization is adopted with a weight decay coefficient of 0.01. The batch size ($N$) is set to 256 and the training process is limited to a maximum of 100 epochs in the pre-training stage and 50 epochs in the fine-tuning stage.
@@ 5.2 Experimental Setup
In each of the three tasks in our self-collection dataset, the experiments are conducted using the Leave-One-Subject-Out (LOSO) strategy. Each subject will be treated as the test subject in turn, which can be considered as the new subject. The remaining subjects constitute the training subject, which can be considered as the existing subjects. To overcome the impact of extreme class imbalance in RSVP tasks, we employ the resampling technique. Specifically, we down-sample the nontarget class to match the sample size of the target class. This down-sampling procedure is only employed in the training set.
We evaluate the effectiveness of TSformer-SA in subject-dependent decoding and subject-independent decoding experiments, respectively. For subject-dependent decoding, the TSformer is initially trained on data from existing subjects in the pre-training stage. Then, the first 4 blocks of data from the new subject are used as the training set to train the subject-specific adapter in the fine-tuning stage, while the remaining blocks serve as the test set. In subject-independent decoding, the TSformer is trained on data from existing subjects and directly tested on the data from new subjects, which ensures that the data from new subjects are not included in the training set.
5.3 Comparison Methods
We compare our proposed model with the following methods:
In their original papers, all of these methods employ subject-dependent decoding settings. For each new subject, the $b$ blocks are designated as the training set, while the remaining blocks are used as the test set. In the practical application of RSVP-BCI, the model is first trained on data collected from new subjects during the preparation procedure. Then, the trained model is utilized to decode the subsequently collected data. To simulate the practical use of RSVP-BCI systems, we utilize the first 4 blocks of each new subject as the training set and the rest blocks serve as the test set in our comparison experiment.
@@ 5.4 Evaluation Metrics
We evaluate model performance using Balanced Accuracy (BA), True Positive Rate (TPR), and False Positive Rate (FPR). The results are presented as mean across all test subjects. The calculation formulas are as follows:
@@
5.5 Statistical Analysis
6 Results and Discussion
6.1 Subject-dependent Experiments
Table 1: Comparisons of BA, TPR, and FPR of TSformer-SA and compared methods in subject-dependent decoding on three tasks (mean).
@@ 6.1 Subject-dependent Experiments
  • 1.
    The asterisks in the table indicate a significant difference between TSformer-SA and the comparison methods by paired t-tests (${}^{\star}$ $p<0.05$, ${}^{\star\star}$ $p<0.01$, ${}^{\star\star\star}$ $p<0.001$). The best results are highlighted in bold and the second-best results are denoted by underline.
@@
6.2 Subject-independent Experiments
Table 2: Comparisons of BA, TPR, and FPR of TSformer and compared methods in subject-independent decoding on three tasks (mean).
@@ 6.2 Subject-independent Experiments
  • 1.
    The asterisks in the table indicate a significant difference between TSformer and the comparison methods by paired t-tests (${}^{\star}$ $p<0.05$, ${}^{\star\star}$ $p<0.01$, ${}^{\star\star\star}$ $p<0.001$). The best results are highlighted in bold and the second-best results are denoted by underline.
@@
6.3 Ablation Study
We conduct ablation studies to evaluate the efficacy of feature extractor, fusion module, and cross-view interaction learning in TSformer-SA. Firstly, to assess the capability of the feature extractor in extracting temporal and spectral features, we only employ the feature extractor followed by a convolution layer and a linear layer for classifying spectrogram images (M1) and EEG temporal signals (M2), respectively. Then, in M3, we employ the feature extractor to extract both temporal and spectral features and these extracted features are then fused in the fusion module for classification. Finally, based on M3, we introduce the cross-view interaction learning to create M4, which is the TSformer-SA. The results for subject-dependent decoding and subject-independent are presented in Table 3 and Table 4, respectively.
Table 3: Ablation study of TSformer-SA in subject-dependent decoding on three tasks (mean).
@@ 6.3 Ablation Study
  • 1.
    The asterisks in the table indicate a significant difference between TSformer-SA and the ablation models by paired t-tests (${}^{\star}$ $p<0.05$, ${}^{\star\star}$ $p<0.01$, ${}^{\star\star\star}$ $p<0.001$). The best results are highlighted in bold and the second-best results are denoted by underline.
  • 2.
    Abbreviations. W: Wavelet view; T: Temporal view; F: Fusion module; C: Cross-view interaction learning.
Table 4: Ablation study of TSformer in subject-independent decoding on three tasks (mean).
@@
  • 1.
    The asterisks in the table indicate a significant difference between TSformer and the ablation models by paired t-tests (${}^{\star}$ $p<0.05$, ${}^{\star\star}$ $p<0.01$, ${}^{\star\star\star}$ $p<0.001$). The best results are highlighted in bold and the second-best results are denoted by underline.
  • 2.
    Abbreviations. W: Wavelet view; T: Temporal view; F: Fusion module; C: Cross-view interaction learning.
@@
The decoding performance of both M1 and M2 indicates that our proposed feature extractor can simultaneously capture the discriminative characteristics of EEG temporal signals as well as spectrogram images. After integrating the features from these two views using the fusion module, M3 exhibits significantly improved BA (all tasks: $p<0.05$) and TPR (all tasks: $p<0.05$) compared to M1 and M2 in three tasks while maintaining a lower FPR in comparison to both models (Task plane: $p=0.38$; Task car and people: $p<0.01$). These results prove the effectiveness of multi-view information for enhancing decoding performance and fusion module in fusing the two view features to obtain more comprehensive fusion features. Finally, our proposed method (M4) incorporates cross-view interaction learning into M3, which significantly enhances the performance on BA (all tasks: $p<0.05$), TPR (all tasks: $p<0.05$), and FPR (Task plane: $p<0.001$; Task car: $p=0.09$; Task people: $p<0.05$) when compared to M3 in three tasks. This proves that our proposed cross-view interaction learning enhances the distinction between fusion features of different classes by improving common feature extraction across two views. Similarly, the results of subject-independent decoding in Table 4 also validate the effectiveness of feature extractor, fusion module, and cross-view interaction learning. By integrating the features from temporal and spectral views with the fusion module, M3 improves the performance of models that use single-view information (M1 and M2) for RSVP decoding. Additionally, our proposed method (M4) significantly outperforms M3 on BA, TPR, and FPR in all three tasks (all tasks: BA, $p<0.05$; TPR, $p<0.05$; FPR, $p<0.01$). Hence, each component proposed in TSformer-SA contributes to enhancing RSVP decoding performance.
@@ 6.4 Efficiency of Subject-specific Adapter
Table 5: Comparisons of BA, TPR, and FPR of TSformer-SA and compared methods using two-stage training strategy on three tasks (mean).
@@
  • 1.
    The asterisks in the table indicate a significant difference between TSformer-SA and the comparison methods by paired t-tests (${}^{\star}$ $p<0.05$, ${}^{\star\star}$ $p<0.01$, ${}^{\star\star\star}$ $p<0.001$). The best results are highlighted in bold and the second-best results are denoted by underline.
@@ 6.5.1 EEG Signals Analysis
We display the grand-average EEG waveforms and spectrogram images of Task plane at Oz and Pz channels in Fig. 5. In Fig. 5 (a) and (b), it is observed that the waveform of the target EEG signals exhibits the negative peak after 200 ms and the positive peak after 300 ms, corresponding to the N200 and P300 components, respectively. Nontarget EEG signals mainly present a harmonic response induced by the 10 Hz picture flicker. The target and nontarget responses of EEG signals are consistent with previous studies polich2007updating ; patel2005characterization .
@@
As shown in Fig. 5 (c) and (d), the spectrogram images of the target EEG signals demonstrate high power within the 0-10 Hz frequency band at around 250 ms, as well as in the interval of 350 ms to 500 ms. Additionally, there is high power in the 0-4 Hz frequency band within the interval of 600 ms to 750 ms. Conversely, since the nontarget EEG signal is a 10 Hz harmonic response, the spectrogram images of nontarget EEG signals exhibit weakly high power which is primarily concentrated in the 5-15 Hz frequency band with 10 Hz as the center frequency at all times. These phenomena are consistent with previous studies makeig2004mining ; kolev1997time . Thus, there are discrepancies between the spectrogram images of target and nontarget EEG signals, which are primarily concentrated within the low-frequency band. These discrepancies can also be utilized for classification.
@@ 6.5.2 Feature Visualization
We employ the t-distributed Stochastic Neighbor Embedding (t-SNE) van2008visualizing to separately project the input data (see Fig. 6 (a)), the output features of feature extractor (see Fig. 6 (b)), the output features of cross-view interaction module (see Fig. 6 (c)), and the output features of fusion module and subject-specific adapter (see Fig. 6 (d)) into a two-dimensional feature space and plotted the scatter plot. We conduct this experiment on the first subject in Task Plane. The TSformer-SA is pre-trained on all other subjects in the Task plane and fine-tuned on the data from the first 4 blocks of the first subject. The remaining blocks are used for visualization. We downsample the nontarget class data in the test set to match the number of samples in the target class for better visualization.
As illustrated in Fig. 6 (a), the input data clusters within the temporal view and spectral view, but both temporal view and spectral view exhibit obvious overlap between target and nontarget classes. After the feature extraction by the feature extractor in each view, the overlap between target and non-target classes within each view is reduced (see Fig. 6 (b)). This indicates that the feature extractor can capture distinctive features from both views. In Fig. 6 (c), following the interaction of two-view features in the cross-view interaction module, the features between temporal and spectral views are aligned and become more closely in the feature space, while the features of target and nontarget classes tend to be linearly separated. This suggests that the cross-view interaction module can effectively extract common representations to reduce the feature gap between the two views. Subsequently, Figure 6 (d) reveals the linear separation between the features output by the fusion module and the subject-specific adapter for the two classes. The fusion module and adapter successfully fuse the features from both views as more distinct fusion features and adapt to the data distribution of new subjects, which achieves superior decoding performance. Thus, our proposed model can effectively realize discriminative feature extraction and feature fusion between the two views for final classification.

### v2

@@ 4.3 Fusion Module
Although temporal tokens and spectral tokens are in different domains, they are naturally aligned along the time dimension. To effectively extract high-level discriminative features of EEG signals, we design an attention-based fusion module, which fuses the temporal and spectral features by taking advantage of the alignment between the two view features and the time-series dependencies of the fused features.
The fusion module is proposed after the cross-view interaction module to fuse temporal and spectral tokens, thereby generating comprehensive fusion features for classification (see Fig. 4). Since temporal tokens and spectral tokens are extracted from EEG temporal signals and spectrogram images during the same time period, they naturally align in the time dimension. As shown in Fig. 4. Initially, we token-wise concatenate the two-view tokens and linearly project them into a $d$-dimensional feature space, each of them containing temporal-spectral information for its corresponding time period. Then we designate these tokens as queries and employ both temporal and spectral tokens as keys and values in a cross-attention layer to obtain fusion tokens ($\boldsymbol{X}_{fus}\in\mathbb{R}^{n\times d}$). At the end of the fusion module, we design a convolution layer with 16 convolution kernels to aggregate all fusion tokens and flatten the output feature into a feature vector as the fusion feature ($\boldsymbol{z}_{fus}$) for final prediction. The convolution kernel size is $(16,d/8)$ with a stride of $(16,d/8)$.
@@ 4.4 Subject-Specific Adapter
Owing to variations in data distribution among different subjects, the pre-trained model needs to transfer the learned general patterns to the data distribution of new subjects for efficient decoding. Moreover, to reduce the preparation time of new subject training data acquisition as much as possible, the amount of new subject data that can be used for fine-tuning the pre-training model is limited. The adapter-based fine-tuning can realize the transfer of pre-trained models to downstream tasks by updating a few parameters, which can avoid overfitting caused by limited data [39]. Thus, we introduce a simple and effective Subject-specific Adapter (SA) to rapidly transfer the knowledge of the pre-trained model learned from existing subjects’ data to data characteristics of new subjects using a small amount of data from the new subject.
@@ 5 Experiments
5.1 Hyperparameters Setting and Implementation Details
The symbols and values of network hyperparameters are summarized in Table 2. We employ the Mexican hat wavelet for the CWT with a scale set to 20, as P300 detection emphasizes low-frequency EEG information. To prevent overfitting, we appropriately reduce the size of the model by setting $d$ as 128 and $h$ as 4.
Table 2: The symbols and values summary of network hyperparameters.
Symbol Definition Value
$C$ Number of EEG channels 64
$T$ Number of EEG sampling points 250
$S$ Size of scale in CWT 20
$t$ Slice length in the slice embedding layer 5
$n$ Length of token sequence 50
$d$ Embedding dimension of all the tokens 128
$h$ Number of heads 4
$\tau$ Temporal parameter of multi-view consistency loss 0.2
In each of the three tasks, each subject will be treated as the test subject in turn, which can be considered as the new subject. The remaining subjects constitute the training subject, which can be considered as the existing subjects. The TSformer-SA is initially trained on data from existing subjects in the pre-training stage using the $\mathcal{L}_{overall}$. Then, the data from the new subject are used as the training set to train the subject-specific adapter in the fine-tuning stage using the $\mathcal{L}_{cross\mbox{-}entropy}$.
We implement our model within the PyTorch framework. In both training stages, we utilize the Adam optimizer [62] for model optimization with a learning rate set to 0.0005, which is reduced by 20% every 40 epochs. Additionally, the L2 regularization is adopted with a weight decay coefficient of 0.01. The batch size ($N$) is set to 256 and the training process is limited to a maximum of 100 epochs in the pre-training stage and 50 epochs in the fine-tuning stage.
(a) Subject-dependent decoding
(b) Subject-independent decoding
(c) Two-stage training
Figure 5: The experimental flow charts of (a) subject-dependent decoding, (b) subject-independent decoding, and (c) two-stage training.
5.2 Comparison Methods
We compare our proposed model with two conventional machine learning methods, five CNN-based methods, and three Transformer-based methods.
For fair comparisons, we resample our EEG signals as described in the comparison methods and modify the channel number to the same as ours. We re-implement the model architecture of the comparison methods exactly and optimize these models using the same optimization techniques and parameter settings as described in the source literature. The comparison methods are as follows:
5.2.1 Conventional machine learning methods
  • HDCA [15]: a linear discrimination method that learns weights on the channel and time window of EEG. We set the time window to 25 sampling points.
  • MDRM [16]: a Riemannian geometry classifier, classifies samples according to the geodesic distance to the center of the category. The method is re-implemented based on the Python package pyriemann.
5.2.2 CNN-based methods
  • MCNN [24]: a three-layer CNN proposed by Manor, we re-implement the network under the description in [24].
  • EEGNet [25]: a CNN-based method with depthwise and separable convolution layers. The implementation refers to https://github.com/vlawhern/arl-eegmodels.
  • LeeNet [43]: a convolutional neural network based on EEGNet, which is suitable for large ERP dataset training. The method is implemented according to [43].
  • PLNet [27]: a CNN-based method using the phase-locked characteristic of ERP signals to extract spatiotemporal features. The method is implemented according to [27].
  • PPNN [28]: a CNN-based method with dilated temporal convolution layers to preserve phase information. The implementation refers to [28].
5.2.3 Transformer-based methods
  • TCN-T [48]: a network architecture based on Temporal Convolution Network and Transformer. The implementation refers to [48].
  • EEG-conformer [47]: a compact convolutional Transformer, which encapsulates local and global features in a unified EEG classification framework. The implementation refers to https://github.com/eeyhsong/EEG-Conformer.
  • CAW-STST [52]: a dual-branch spatio-temporal-spectral Transformer with a channel attention weighting module. The implementation refers to https://github.com/ljbuaa/VisualDecoding.
@@ 5.3 Experimental Setup
In section 3, we conduct three distinct RSVP experiments for three different application scenarios. The subjects in the three experiments are not repeated. Consequently, the Task plane, Task car, and Task people can be viewed as three independent RSVP datasets. When evaluating the model’s performance, we conduct all experiments on three datasets separately.
To assess our model’s performance, we first compare it with subject-dependent decoding methods. Next, we evaluate TSformer’s capability to extract task-related features from existing subjects and generalize them to new subjects by comparing the subject-independent decoding performance of TSformer and comparison methods. To validate the effectiveness of the proposed subject-specific adapter, we compare our method with comparison methods using the same two-stage training setting. The process of these three experimental settings are as follows:
  • 1)
    Subject-dependent decoding (see Fig 5 (a)) requires the training set and test set from the same subject collected in the same RSVP task. In each of the three tasks, for each subject, the data of the first $b$ ($b=1,2,3,4$) blocks are designated as the training set, while the remaining blocks are used as the test set. This setting simulates the practical use of RSVP-BCI systems.
  • 2)
    Subject-independent decoding (see Fig 5 (b)) requires the training set and test set from different subjects collected in the same RSVP task. In each of the three tasks, the experiments are conducted using the Leave-One-Subject-Out (LOSO) settings. In each task, each subject will be treated as the test set in turn and the remaining subjects constitute the training set.
  • 3)
    Two-stage training strategy (see Fig 5 (c)) involves the pre-training stage and fine-tuning stage. For each task, every subject takes a turn as the test subject. First, the model is pre-trained using the data from other subjects in the same RSVP task in the pre-training stage. Then, the pre-trained model is fine-tuned using the data from the first b blocks of the test subjects. Finally, the model’s performance is evaluated on the remaining data from the test subjects.
Moreover, to overcome the impact of extreme class imbalance in RSVP tasks, we employ the resampling technique. Specifically, we down-sample the nontarget class to match the sample size of the target class. This down-sampling procedure is only employed in the training set.
@@ 5.4 Evaluation Metrics
We evaluate model performance using Balanced Accuracy (BA), True Positive Rate (TPR), and False Positive Rate (FPR). All experiments are conducted separately on each of the three datasets, so the average of BA, TPR, and FPR for each test subject in the three tasks are calculated separately. The calculation formulas are as follows:
@@
Table 3: Comparisons of BA, TPR, and FPR of TSformer-SA and compared methods in subject-dependent decoding on three test tasks (mean).
@@
  • The asterisks in the table indicate a significant difference between TSformer-SA and the comparison methods by paired t-tests ( $p<0.05$, ⋆⋆ $p<0.01$, ⋆⋆⋆ $p<0.001$). The best results are highlighted in bold and the second-best results are denoted by underline.
5.5 Statistical Analysis
6 Results and Discussion
6.1 Subject-Dependent Experiments
@@
Table 4: Comparisons of BA, TPR, and FPR of TSformer and compared methods in subject-independent decoding on three test tasks (mean).
@@
  • The asterisks in the table indicate a significant difference between TSformer and the comparison methods by paired t-tests ( $p<0.05$, ⋆⋆ $p<0.01$, ⋆⋆⋆ $p<0.001$). The best results are highlighted in bold and the second-best results are denoted by underline.
6.2 Subject-Independent Experiments
@@
Table 5: Ablation study of TSformer-SA in subject-dependent decoding on three test tasks (mean).
@@
  • The asterisks in the table indicate a significant difference between TSformer-SA and the ablation models by paired t-tests ( $p<0.05$, ⋆⋆ $p<0.01$, ⋆⋆⋆ $p<0.001$). The best results are highlighted in bold and the second-best results are denoted by underline.
  • Abbreviations. W: Wavelet view; T: Temporal view; F: Fusion module; C: Cross-view interaction module.
Table 6: Ablation study of TSformer in subject-independent decoding on three test tasks (mean).
@@
  • The asterisks in the table indicate a significant difference between TSformer and the ablation models by paired t-tests ( $p<0.05$, ⋆⋆ $p<0.01$, ⋆⋆⋆ $p<0.001$). The best results are highlighted in bold and the second-best results are denoted by underline.
  • Abbreviations. W: Wavelet view; T: Temporal view; F: Fusion module; C: Cross-view interaction module.
6.3 Ablation Study
We conduct ablation studies to evaluate the efficacy of the feature extractor, fusion module, and cross-view interaction module in TSformer-SA. Firstly, to assess the capability of the feature extractor in extracting temporal and spectral features, we only employ the feature extractor followed by a convolution layer and a linear layer for classifying spectrogram images (M1) and EEG temporal signals (M2), respectively. Then, in M3, we employ the feature extractor to extract both temporal and spectral features and these extracted features are then fused in the fusion module for classification. Finally, based on M3, we introduce the cross-view interaction module to create M4, which is the TSformer-SA. The results for subject-dependent decoding and subject-independent are presented in Table 5 and Table 6, respectively.
@@ 6.3 Ablation Study
The decoding performance of both M1 and M2 indicates that our proposed feature extractor can simultaneously capture the discriminative characteristics of EEG temporal signals as well as spectrogram images. After integrating the features from these two views using the fusion module, M3 exhibits significantly improved BA (all tasks: $p<0.05$) and TPR (all tasks: $p<0.05$) compared to M1 and M2 in three tasks while maintaining a lower FPR in comparison to both models (Task plane: $p=0.38$; Task car and people: $p<0.01$). These results prove the effectiveness of multi-view information for enhancing decoding performance and fusion module in fusing the two view features to obtain more comprehensive fusion features. Finally, our proposed method (M4) incorporates cross-view interaction module into M3, which significantly enhances the performance on BA (all tasks: $p<0.05$), TPR (all tasks: $p<0.05$), and FPR (Task plane: $p<0.001$; Task car: $p=0.09$; Task people: $p<0.05$) when compared to M3 in three tasks. This proves that our proposed cross-view interaction module enhances the distinction between fusion features of different classes by improving common feature extraction across two views. Similarly, the results of subject-independent decoding in Table 6 also validate the effectiveness of the feature extractor, fusion module, and cross-view interaction module. By integrating the features from temporal and spectral views with the fusion module, M3 improves the performance of models that use single-view information (M1 and M2) for RSVP decoding. Additionally, our proposed method (M4) significantly outperforms M3 on BA, TPR, and FPR in all three tasks (all tasks: BA, $p<0.05$; TPR, $p<0.05$; FPR, $p<0.01$). Hence, each component proposed in TSformer-SA contributes to enhancing RSVP decoding performance.
@@ 6.4 Efficiency of Subject-Specific Adapter
Table 7: Comparisons of BA, TPR, and FPR of TSformer-SA and compared methods using two-stage training strategy on three test tasks (mean).
@@
  • The asterisks in the table indicate a significant difference between TSformer-SA and the comparison methods by paired t-tests ( $p<0.05$, ⋆⋆ $p<0.01$, ⋆⋆⋆ $p<0.001$). The best results are highlighted in bold and the second-best results are denoted by underline.
@@ 6.5.1 EEG Signals Analysis
We display the grand-average EEG waveforms and spectrogram images of Task plane at Oz and Pz channels in Fig. 6. In Fig. 6 (a) and (b), it is observed that the waveform of the target EEG signals exhibits the negative peak after 200 ms and the positive peak after 300 ms, corresponding to the N200 and P300 components, respectively. Nontarget EEG signals mainly present a harmonic response induced by the 10 Hz picture flicker. The target and nontarget responses of EEG signals are consistent with previous studies [14, 63].
As shown in Fig. 6 (c) and (d), the spectrogram images of the target EEG signals demonstrate high power within the 0-10 Hz frequency band at around 250 ms, as well as in the interval of 350 ms to 500 ms. Additionally, there is high power in the 0-4 Hz frequency band within the interval of 600 ms to 750 ms. Conversely, since the nontarget EEG signal is a 10 Hz harmonic response, the spectrogram images of nontarget EEG signals exhibit weakly high power which is primarily concentrated in the 5-15 Hz frequency band with 10 Hz as the center frequency at all times. These phenomena are consistent with previous studies [18, 64]. Thus, there are discrepancies between the spectrogram images of target and nontarget EEG signals, which are primarily concentrated within the low-frequency band. These discrepancies can also be utilized for classification.
@@ 6.5.2 Feature Visualization
We employ the t-distributed Stochastic Neighbor Embedding (t-SNE) [65] to separately project the input data (see Fig. 7 (a)), the output features of feature extractor (see Fig. 7 (b)), the output features of cross-view interaction module (see Fig. 7 (c)), and the output features of fusion module and subject-specific adapter (see Fig. 7 (d)) into a two-dimensional feature space and plotted the scatter plot. We conduct this experiment on the first subject in Task Plane. The TSformer-SA is pre-trained on all other subjects in the Task plane and fine-tuned on the data from the first 4 blocks of the first subject. The remaining blocks are used for visualization. We downsample the nontarget class data in the test set to match the number of samples in the target class for better visualization.
As illustrated in Fig. 7 (a), the input data clusters within the temporal view and spectral view, but both temporal view and spectral view exhibit obvious overlap between target and nontarget classes. After the feature extraction by the feature extractor in each view, the overlap between target and non-target classes within each view is reduced (see Fig. 7 (b)). This indicates that the feature extractor can capture distinctive features from both views. In Fig. 7 (c), following the interaction of two-view features in the cross-view interaction module, the features between temporal and spectral views are aligned and become more closely in the feature space, while the features of target and nontarget classes tend to be linearly separated. This suggests that the cross-view interaction module can effectively extract common representations to reduce the feature gap between the two views. Subsequently, Fig. 7 (d) reveals the linear separation between the output features by the fusion module and the subject-specific adapter for the two classes. The fusion module and adapter successfully fuse the features from both views as more distinct fusion features and adapt to the data distribution of new subjects, which achieves superior decoding performance. Thus, our proposed model can effectively realize discriminative feature extraction and feature fusion between the two views for final classification.
6.5.3 Subject Task-Related Feature Extraction
In Fig. 8 (a), it can be observed that due to variations between different subjects and classes, EEG temporal signals exhibit very low similarity among subjects, with only weak similarities in the target data from different subjects. After feature extraction, the similarity of features among subjects increases (see Fig. 8 (b)). However, there are also notable similarities among features from different classes, which may increase FPR in classification. Figure 8 (c) illustrates that the similarity between features from different subjects within the same class is significantly enhanced compared to Fig. 8 (b) and is much higher than the similarity between features from different classes. This improvement strengthens the TPR while reducing FPR in classification. These results demonstrate the model’s ability to effectively extract common task-related features across subjects through pre-training on data from various subjects.